# Internal Consistency and Self-Feedback in Large Language Models: A Survey

### v2

@@ II-C Status Quo of LLM Consistency
This section briefly discusses the current performance of LLMs in terms of consistency. As indicated at the beginning of the survey, GPT-4o’s various responses to the same question (see Fig. 1) already demonstrate that even relatively powerful language models still exhibit low consistency.
There are also many works that either actively or inadvertently explore the Internal Consistency of LLMs. The well-known Self-Consistency [2] explores the use of the majority voting strategy, where the LLM generates multiple responses and selects the most voted one as the final response. Their experiments showed that on the reasoning benchmark GSM8K [34], this method increased the answer accuracy from 56.5% to 74.4%. This implies that many responses may be random and do not represent a consistent response.
In terms of hallucination alleviation, Mündle et al. [21] proposed the Self-Contradict strategy, which attempts to generate different samples to identify self-contradictory content and then eliminate these contradictions to reduce hallucinations. Their experiment showed that GPT-4, ChatGPT, Llama2-70B-Chat, and Vicuna-13B were able to induce self-contradictions at rates of 15.7%, 17.7%, 19.0%, and 22.9%, respectively.
Empirical tests show LLMs’ direct responses are inconsistent, even when they know the correct answer.
Besides response consistency, some works have explored whether models can consistently express what they know and do not know, a capability referred to as Self-Knowledge. For example, Yin et al. [4] and Cheng et al. [35] created datasets consisting of questions that models cannot answer to test whether the models can refuse to answer these questions. Their research showed that models exhibit low consistency in refusing “I Don’t Know” (IDK) questions, with room for improvement compared to humans.
@@ II-D Sources of Low Internal Consistency
Why do models exhibit low consistency? Many scholars have conducted in-depth research on this phenomenon, exploring the causes from various angles such as prompt engineering, the decoding process, and the attention mechanism.
@@
Some studies explore the impact of slight variations in prompts on response consistency. The structures of prompt may cause low consistency. Xie et al. [36] designed different CoT prompts and discovered that under different CoT’s guidance, the latent state’s distance between the intermediate and final layers of LLMs varied significantly. Simply put, some prompt designs result in low consistency between different latent layers, with significant differences in the probability distribution of the next token prediction. Liu et al. [37] tested the accuracy of LLMs’ responses by placing text containing answers in different positions within the prompt, discovering a “lost-in-the-middle” phenomenon. Models tend to focus on the content at the beginning and end of prompts, leading to inconsistent responses to differently structured prompts. Similarly, Liu et al. [38] found that hallucinations emerge when LLMs dealing with long context. They analyzed that this is caused by the soft attention mechanism, where attention weights become overly dispersed as sequence length increases, leading to poor consistency in reasoning paths.
@@
In addition to prompt engineering explorations, some researchers have directly investigated drawbacks within the LLM architecture. Yang et al. [33] studied whether models perform intermediate reasoning within latent states when answering questions and whether enhancing this intermediate reasoning signal strength could improve answer accuracy. Taking the example of the model answering “In what year did Plato’s teacher die? [^7] Plato’s teacher is Socrates, who died in 399 BC.”, this work specifically investigated whether the model infers the intermediate entity “Socrates” in the latent layer and whether increasing the weight of the latent state corresponding to the intermediate entity would make the model answer more accurately. The experimental results showed that models do possess latent reasoning capabilities, but these are weak. On the one hand, the signal strength of the intermediate entity is weak. On the other hand, enhancing this signal strength did not significantly improve the LLM’s response. This indicates deficiencies in current LLM architectures in performing and expressing latent reasoning. This means that when predicting the next token, the model may make near-random predictions due to failed intermediate entity reasoning. Additionally, Zhang et al. [39] argued that models could appear hallucinations due to the “snowball effect”. The full attention mechanism makes LLMs overly confident in their outputs, leading to compounding errors if an initial reasoning mistake occurs. Consequently, model’s responses may become inconsistent with the knowledge it has learned.
@@
Furthermore, some hypotheses and viewpoints reveal potential reasons for the low internal consistency of current LLMs. Bender et al. [40] proposed that large language models might be “stochastic parrots”, learning rules and patterns from training data rather than truly understanding the grammar and semantics of natural language. This inherent randomness in generation reflects a form of internal inconsistency in the model. Ma et al. [41] proposed the Principle of Self-Consistency for intelligent agents, aiming to find a coherent model that minimizes internal differences between observed and regenerated data. They found many factors that could affect internal consistency, such as mode collapse [^8] Mode collapse: A generative model starts producing very similar or repetitive outputs during training, failing to capture the diversity of the data., neural collapse [^9] Neural collapse: The model learns the simplest representation to map input to output, without capturing the complex logic within the data., and over-fitting or under-fitting caused by overly high or low dimensional feature spaces.
In conclusion, both theoretical and experimental findings indicate that model architecture, training processes, and user’s queries can all contribute to low internal consistency. Understanding these causes can help researchers better address this issue and improve model performance.
@@ II-E How to Understand Internal Consistency?
This section adopts a broader perspective, an alignment perspective, to help readers understand Internal Consistency. If there is Internal Consistency, there must also be corresponding External Consistency. Internal Consistency focuses on whether the model can align with itself during expression. External Consistency includes the alignment between the pre-training dataset and the pre-training model parameters, the alignment between the pre-trained model and the chat model, and the alignment between the chat model and the model subjected to RLHF (Reinforcement Learning with Human Feedback), among others. These different alignments can be illustrated in Fig. 6. Each stage of alignment plays a unique role. Internal Consistency is crucial for AI safety. Kadavath et al. [42] mention the significant value of Internal Consistency:
Figure 6: Various Alignments Involved in the LLM development
  • Truthfulness. LLMs should provide factually accurate information, including finding, using, and evaluating source materials correctly.
  • Calibration. LLMs’ probabilistic predictions should correspond with frequencies of occurrence.
  • Self-Knowledge. LLMs should know what they know and make accurate predictions about their own behavior.
  • Explainability. LLMs should reveal their “thinking” completely and faithfully.
  • Non-deceptiveness. LLMs should be ensured not to lie, even when human preference might encourage systematic mistakes or provide rewards for pleasant misconceptions.
In addition to these significant values in AI safety, Internal Consistency also affects the robustness and reliability of AI systems. For instance, the current inability of models to reason consistently directly impacts an AI agent’s understanding of goals, leading to incorrect operations by the agent [43].
@@ III-A Formulation
Narrowly speaking, Self-Feedback refers to the method of improving a model’s own Internal Consistency through its feedback, where “own” refers to a specific model entity or a specific response. Broadly speaking, “own” can be extended to other models. For example, multiple different models can improve their capabilities through feedback generated from debates among them, which is a more generalized interpretation of Self-Feedback.
Based on the above descriptive definition, we can formalize the process of Self-Feedback. For a given model $\mathcal{M}$, query $\bm{x}$, and a sampling set $\mathcal{Y}$ obtained under a certain expression, Self-Evaluate [^10] A small number of methods use other models $\text{SelfEvaluate}_{\mathcal{N}}(\mathcal{Y})$ or even external tools $\text{SelfEvaluate}_{\text{tool}}(\mathcal{Y})$ during Self-Evaluate. is first performed to obtain feedback $f$:
@@ III-B Taxonomy
The Self-Feedback methods introduced in this paper can be categorized from two different perspectives, as shown in the middle and bottom parts of Fig. 3.
The first perspective classifies based on the components of Self-Feedback, as depicted in the middle part of Fig. 3. In Self-Feedback, the initial step is Self-Evaluation. The model can evaluate its own response, token probability distribution, or the latent states. Through Self-Evaluation, different consistency signals can be obtained: scalar signals (e.g., the confidence level of the output), textual signals (e.g., the model’s critique of its own output), external signals (e.g., results from a Python interpreter), and comparative signals (e.g., different token probability distributions induced by two different prompts). These signals are then used for downstream Self-Update, with three typical update strategies. The model can use the signals to prompt itself for direct result modification, identify better outputs for re-tuning itself through Self-Evaluation, or even re-tune a student model in the domain of knowledge distillation using better outputs. Theoretically, this yields a total of $3\times 4\times 3=36$ method combinations. Obviously, the drawback of this taxonomy is its excessive granularity and lack of task-oriented perspective, which make it difficult for researchers to design new solutions for specific task.
To avoid the shortcoming, we primarily adopt the second perspective, categorizing based on the tasks and lines of work, as illustrated at the bottom of Fig. 3. Firstly, the major challenge for LLMs is low Internal Consistency. To address this, it is essential to clarify what the intermediate consistency signals are and how they are acquired. In Section IV, we summarize six important lines of work.
In terms of specific manifestations, low Internal Consistency includes two major issues: weak reasoning ability in QA tasks and hallucinations in open-ended generation. Therefore, in Section V, we summarize three lines of work for Reasoning Elevation. And in Section VI, we summarize four lines of work for Hallucination Alleviation. Table IV summarizes these seven lines of work. Additionally, some tasks do not aim to improve consistency but still utilize the Self-Feedback framework. For comprehensiveness, we also briefly summarize these other tasks in Section VII.
@@
1 No Majority Voting, Value Function Best Selection Self-Consistency [2], ToT [44], GoT [45]
V-B: Refining with Responses Response Textual 1 or 2 Half Sampling Best Selection, Model Tuning Self-Improve [46], ConCoRD [47], LEMA [48]
V-C: Multi-Agent Collaboration Response Textual, Scalar $\geq 2$ Rare Negotiation Answer Aggregation FORD [49], MACNet [50], REFINER [51]
VI-A: Refining the Response Iteratively Response Textual, External 1 Few Model Generate Critique Model Generate Refinement Self-Refine [8], Reflexion [52], Self-Correct [9]
@@ IV Task: Consistency Signal Acquisition
Consistency signal acquisition refers to evaluating the consistency of expressions after obtaining the sampling set $\mathcal{Y}$ from the language model $\mathcal{M}$ for the query $\bm{x}$. The evaluated signal can help the model update its expressions or parameters, thereby improving the model’s Internal Consistency. Therefore, consistency signal acquisition is a pivotal task within the Self-Feedback framework. These methods either require access only to the model’s output contents, to the logits, or the latent states of the model. Depending on the depth of access required by different methods, the approaches mentioned in this section are categorized as black-box (accessing only the model’s output contents), gray-box (also accessing logits), and white-box (also accessing the model’s latent states). Numerous explorations have been undertaken in this task. These include:
@@
The names of the first three lines of work mentioned above separate research topics that are actually quite similar. They all provide scalar feedback for LLM responses, and some works even mix the keywords from these three lines, such as [58, 59, 60]. The differences among these three lines mainly lie in the slight distinctions in their downstream tasks. Estimating the uncertainty and confidence of model expressions is essentially two sides of the same coin, as both calculate the model’s certainty level to obtain a scalar within the range of $[0,1]$. This approach is often used to optimize the model’s reasoning ability, selecting a better reasoning path based on the uncertainty rate or confidence rate. Moreover, hallucination detection determines the presence of hallucinations (choosing from $\{0,1\}$) through various methods, which is evidently more used in the task of hallucination alleviation.
@@ IV-A Uncertainty Estimation
Due to the black-box nature of the deep learning, uncertainty estimation has always been an important topic. Uncertainty estimation refers to estimating the data uncertainty, model uncertainty, and distributional uncertainty involved in the neural networks [61]. For uncertainty estimation in the NLP field, Hu et al. [25] conducted a detailed survey. Interested readers can refer to this article for further understanding. Here we briefly introduce the uncertainty modeling proposed in their work.
The purpose of uncertainty modeling is to identify the sources of uncertainty that cause the model to generate uncertain results and to systematically understand the uncertainties present in the model. Hu et al. [25] categorize and explain the sources and modeling methods of uncertainty from the perspectives of models, outputs, and distributions. 1) Calibration Confidence-based Methods: This approach aims to correct the reliability of the uncertainty estimates provided by the model. The basic idea is to compare the accuracy of predicted probabilities with actual probabilities. 2) Sampling-based Methods: This approach models the variability of multiple expressions provided by the model, allowing us to observe the arising uncertainties. This method is also the focus of our article. 3) Distribution-based Methods: This approach starts with the training dataset, pre-judging the distribution characteristics of the dataset and directly calculating the model’s inherent uncertainty by constructing specific distribution functions.
@@ IV-B Confidence Estimation
Literally, confidence is just the antonym of uncertainty, and their evaluation objectives are the opposite. Uncertainty tends to provide insights into model interpretability, while confidence focuses on providing reliability scores, thereby directly enhancing user trust and engagement with the responses.
In this line of work, Self-Evaluation is the core method [^11] The Self-Evaluation here denotes the method proposed in [42], not the Self-Evaluation in Self-Feedback framework. To distinguish between the two, citation markers will be appended when referring to the method.. The concept of Self-Evaluation was first proposed in [42], where the goal is for the model to express its level of confidence using its own knowledge and reasoning. As shown in Fig. 7, the Self-Evaluation method simply asks the AI: Is the proposed answer True or False? Then, the confidence score, P(True), is extracted from the model’s logits.
@@
Besides naively asking the model whether it thinks the proposed answer is correct, some works have proposed other frameworks. For instance, BSDetector [64] is a confidence estimation framework suitable for both black-box and white-box models. It combines the consistency of multiple outputs sampled from the model with model’s own reflection on its output, weighting these scores to obtain the confidence scores. Another example, TrustScore [65] is a reference-free confidence estimation framework using behavior consistency. It generates distractors based on entity information rules from Wikipedia, asks the LLM multiple times, and checks if it consistently chooses its own generated answer.
@@ IV-C Hallucination Detection
The Alibaba team proposed INSIDE [67] for deeper exploration. They sampled the latent expressions. They extracted the vectors from the intermediate layers corresponding to the last token of the sampled sentences and calculated the covariance matrix of these vectors. Since the eigenvalue of the covariance matrix represents data variability, they used this value as a measure of hallucination. Intuitively, if the variability is small, it indicates that the sentence vectors are not far apart from each other, and the likelihood of hallucination is low.
The above works generally employ a single LLM. Additionally, some methods utilize multiple agents to detect hallucination. For example, Cross Examination [68] uses two LLMs, an Examinee, and an Examiner, employing a cross-examination approach to determine factual errors. The Examinee generates responses, while the Examiner iteratively asks questions related to the initial response. The Examinee answers these questions for several rounds, then examiner determines whether there are factual errors.
@@ IV-D Verbal Critiquing
The aforementioned methods can only provide scalar signals for downstream Self-Update, lacking flexibility. Inspired by the idea that “all tasks are generation tasks” [69, 70], many works have proposed allowing LLMs to generate more semantically rich textual signals. There include:
@@ IV-E Contrastive Optimization
Here are some examples of scoring functions at different layers. At the latent layer, in order to find attention heads with a stronger preference for truthfulness, Li et al. [5] trained a probe to evaluate the attention head’s ability of answering questions truthfully. At the decoding layer, Self-Evaluation [42] can be used to evaluate the reasoning paths during beam search, comparing scores to choose a better decoding direction [74]. At the response layer, the well-known Self-Consistency [2] strategy implicitly relies on comparisons between different responses. A variant, Soft Self-Consistency [75], calculates the joint probability of tokens for each response as the scoring function.

### v3

@@ II-C Status Quo of LLM Consistency
As indicated at the beginning of the survey, GPT-4o’s various responses to the same question (see Fig. 1) already demonstrate that even relatively powerful LLMs still exhibit low consistency. This section examines the current state of LLM consistency from two perspectives.
LLMs often provide inconsistent responses, even when they know the correct answer. The well-known Self-Consistency [2] explores the use of the majority voting strategy, where the LLM generates multiple responses and selects the most voted one as the final response. Their experiments showed that on the reasoning benchmark GSM8K [27], this method increased the answer accuracy by about 17.9%. This implies that many initial responses do not represent a consistent answer. In terms of hallucination alleviation, Mündle et al. [18] proposed the Self-Contradict strategy, which attempts to generate different samples to identify self-contradictory content and then eliminate these contradictions to reduce hallucinations. Their experiment showed that even GPT-4 was able to induce self-contradictions at rates of 15.7%.
LLMs are inconsistent in expressing what they know and do not know, i.e., they lack Self-Knowledge. For example, Yin et al. [4] and Cheng et al. [28] created datasets consisting of questions that models cannot answer to test whether the models can refuse to answer these questions. Their research showed that models exhibit low consistency in refusing “I Don’t Know” (IDK) questions, with room for improvement compared to humans.
@@ II-D Sources of Low Internal Consistency
Why models exhibit low internal consistency. Here we present some relevant explorations. Understanding these causes can help researchers better improve model performance.
@@
Great sensitivity to specific prompts. Xie et al. [29] found that different CoT prompts led to significant differences in latent state distances between intermediate and final layers, affecting consistency. Liu et al. [30] observed a “lost-in-the-middle” phenomenon, where models inconsistently respond to prompts based on the position of answers within the long context. Liu et al. [31] further analyzed hallucinations within long contexts. They analyzed that this is caused by the soft attention mechanism, where attention weights become overly dispersed as sequence length increases, leading to poor consistency in reasoning paths.
@@
Deficiencies of reasoning. Yang et al. [26] investigated whether models use intermediate latent reasoning for answering questions and if strengthening this reasoning could boost accuracy. Their findings revealed that while models do have latent reasoning abilities, these are weak. Enhancing the signal strength of intermediate entities did not significantly improve the model’s responses, suggesting current LLM architectures struggle with latent reasoning and may make near-random predictions due to insufficient latent reasoning. Additionally, Zhang et al. [32] argued that models could hallucinate due to the “snowball effect”. The full attention mechanism makes LLMs overly confident in their outputs, leading to compounding errors if an initial reasoning mistake occurs. Consequently, model’s responses may become inconsistent with the knowledge it has learned.
@@
Theoretical hypotheses. Bender et al. [33] proposed that LLMs might be “stochastic parrots”, learning rules and patterns from training data rather than truly understanding the grammar and semantics of natural language. This inherent randomness in generation reflects a form of internal inconsistency in the model. Ma et al. [34] proposed the Principle of Self-Consistency for intelligent agents, aiming to find a coherent model that minimizes internal differences between observed and regenerated data. They found many factors that could affect internal consistency, such as mode collapse [^4] Mode collapse: A generative model starts producing very similar or repetitive outputs during training, failing to capture the diversity of the data., neural collapse [^5] Neural collapse: The model learns the simplest representation to map input to output, without capturing the complex logic within the data., and over-fitting or under-fitting caused by overly high or low dimensional feature spaces.
@@ II-E How to Understand Internal Consistency?
If there is internal consistency, there must also be corresponding external consistency as illustrated in Fig. 6. Each stage of alignment plays a unique role. Among these alignments, internal consistency is crucial for AI reliability [35, 36]:
  • Truthfulness. LLMs provide factually accurate information, including finding, using, and evaluating source materials correctly.
  • Calibration. LLMs’ probabilistic predictions correspond with frequencies of occurrence.
  • Self-Knowledge. LLMs know what they know and make accurate predictions about their own behavior.
  • Explainability. LLMs reveal their “thinking” completely and faithfully.
  • Non-deceptiveness. LLMs are ensured not to lie, even when human preference encourages systematic mistakes or provides rewards for pleasant misconceptions.
@@ III-A Formulation
Narrowly speaking, Self-Feedback refers to the method of improving a model’s own internal consistency through its feedback, where “own” refers to a specific model instance or a specific response. Broadly speaking, “own” can be extended to other models. For example, multiple different models can improve their capabilities through feedback generated from debates among them, which is a more generalized interpretation of Self-Feedback.
Based on the above descriptive definition, we can formalize the process of Self-Feedback. For a given model $\mathcal{M}$, query $\bm{x}$, and a sampling set $\mathcal{Y}$ obtained under a certain expression type, Self-Evaluate [^6] A small number of methods use other models $\text{SelfEvaluate}_{\mathcal{N}}(\mathcal{Y})$ or even external tools $\text{SelfEvaluate}_{\text{tool}}(\mathcal{Y})$ during Self-Evaluate. is first performed to obtain feedback $f$:
@@ III-B Taxonomy
Self-Feedback centers on SelfEvaluate, SelfUpdate, and the feedback signal $f$. Rather than fragmenting the survey by these elements, we classify the papers we read by tasks and lines of work, as shown in Fig. 3. The four key tasks are:
  • Section IV (Consistency Signal Acquisition) summarizes methods for obtaining the feedback signal $f$. We consider this task important because many Self-Feedback methods overlook this dimension. For instance, the feedback signal in Self-Consistency [2] should be classified under scalar-based Consistency Estimation methods.
  • Section V (Reasoning Elevation) is one of the key focuses of this paper. We have discussed the distinctions and connections between reasoning and hallucination in Section I-A. To clarify, the primary focus here is on Self-Feedback methods aimed at QA tasks.
  • Section VI (Hallucination Alleviation) is another critical focus of this paper. Here, we concentrate on Self-Feedback methods targeted at open-ended generation tasks. Note: We also provide Table IV to share specific lines of work related to Reasoning Elevation and Hallucination Alleviation.
  • Section VII (Others) briefly covers Self-Feedback methods applied to tasks beyond Reasoning Elevation and Hallucination Alleviation, such as knowledge distillation and embedding generation.
@@
1 No Majority Voting, Value Function Best Selection Self-Consistency [2], ToT [37], GoT [38], Quiet-STaR [39]
V-B: Refining with Responses Response Textual 1 or 2 Half Sampling Best Selection, Model Tuning Self-Improve [40], ConCoRD [41], LEMA [42], Mistake Tuning [43]
V-C: Multi-Agent Collaboration Response Textual, Scalar $\geq 2$ Rare Negotiation Answer Aggregation FORD [44], MACNet [45], REFINER [46], Multi-Agent Debate [47]
VI-A: Refining the Response Iteratively Response Textual, External 1 Few Model Generate Critique Model Generate Refinement Self-Refine [8], Reflexion [48], Self-Correct [9], Self-Debug [49]
@@ IV Task: Consistency Signal Acquisition
Consistency signal acquisition refers to evaluating the consistency of expressions after obtaining the sampling set $\mathcal{Y}$. The evaluated signal can help the model update its expressions or parameters, thereby improving the model’s internal consistency. Therefore, it is a pivotal task within the Self-Feedback framework. These methods either require access only to the model’s output contents, to the logits, or to the latent states of the model. Depending on the depth of access required by different methods, the approaches mentioned in this section are categorized as black-box (accessing only the model’s output contents), gray-box (also accessing logits), and white-box (also accessing the model’s latent states). Numerous explorations have been undertaken in this task. These include:
@@
The first three lines are actually quite similar. They all provide scalar feedback for LLM responses, and some works even mix the keywords from these three lines, such as [55, 56, 57]. The main difference lies in their downstream tasks. Estimating uncertainty and confidence are two sides of the same coin, both assessing the model’s certainty on a $[0,1]$ scale to optimize reasoning. While hallucination detection identifies hallucinations from $\{0,1\}$, primarily aimed at alleviating hallucinations.
@@ IV-A Uncertainty Estimation
Uncertainty estimation refers to estimating the data uncertainty, model uncertainty, and distributional uncertainty involved in the neural networks [58].
For uncertainty estimation in the NLP field, Hu et al. [24] conducted a detailed survey. They categorize sources and modeling methods of uncertainty into three approaches: 1) Calibration Confidence-based Methods: This approach compares the accuracy of predicted probabilities with actual probabilities. 2) Sampling-based Methods: This approach models the variability of multiple expressions provided by the model, allowing us to observe the arising uncertainties. This method is also the focus of our article. 3) Distribution-based Methods: This approach evaluates inherent uncertainty by analyzing the dataset’s distribution characteristics.
@@ IV-B Confidence Estimation
Confidence is the opposite of uncertainty, focusing on reliability scores to enhance user trust.
In this line of work, Self-Evaluation is the core method [^7] The Self-Evaluation [36] here denotes a method, not the Self-Evaluation module in Self-Feedback framework. To distinguish between the two, a citation marker will be appended when referring to the method.. The concept of Self-Evaluation was first proposed in [36], where the goal is for the model to express its level of confidence using its own knowledge and reasoning. As shown in Fig. 7, the Self-Evaluation method simply asks the model: Is the proposed answer True or False? Then, the confidence score, P(True), is extracted from the model’s logits.
@@
Besides naively asking the model whether it thinks the proposed answer is correct, some works have proposed other frameworks. For instance, BSDetector [61] is a confidence estimation framework suitable for both black-box and white-box models. It combines the consistency of multiple outputs sampled from the model with the model’s own reflection on its output, weighting these scores to obtain the confidence scores. Another example, TrustScore [62] is a reference-free confidence estimation framework using behavior consistency. It generates distractors based on entity information rules from Wikipedia, asks the LLM multiple times, and checks if it consistently chooses its own generated answer.
@@ IV-C Hallucination Detection
Beyond the extensive methods of SelfCheckGPT, there are other interesting approaches as well. The Alibaba team proposed INSIDE [64] for deeper exploration. They sampled latent vectors from the intermediate layers and calculated the covariance matrix of these vectors. Since the eigenvalue of the covariance matrix represents data variability, they used this value as a measure of hallucination. Additionally, some methods utilize multiple agents to detect hallucinations. For example, Cross Examination [65] employs two LLMs, an Examinee and an Examiner, using a cross-examination approach to determine factual errors.
@@ IV-D Verbal Critiquing
Inspired by the idea that “all tasks are generation tasks” [66, 67], many works have proposed allowing LLMs to generate more semantically rich textual signals. These include:
@@ IV-E Contrastive Optimization
At the latent layer, in order to find attention heads with a stronger preference for truthfulness, Li et al. [5] trained a probe to evaluate the attention head’s ability to answer questions truthfully. At the decoding layer, Self-Evaluation [36] can be used to evaluate the reasoning paths during beam search, comparing scores to choose a better decoding direction [70]. At the response layer, Self-Consistency [2] strategy implicitly relies on comparisons between different responses. A variant, Soft Self-Consistency [71], calculates the joint probability of tokens for each response as the scoring function.